Find the difference:
7-(-5)
A.-12
B.2
C.12
D.-2

Answer by jim_thompson5910(35256)

7-(-5) = 7+5 = 12

So the answer is C) 12


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim